### v2

# Simultaneous Calibration of Noise Covariance and Kinematics for State Estimation of Legged Robots via Bi-level Optimization

Abstract
Accurate state estimation is critical for legged and aerial robots operating in dynamic, uncertain environments. A key challenge lies in specifying process and measurement noise covariances, which are typically unknown or manually tuned. In this work, we introduce a bi-level optimization framework that jointly calibrates covariance matrices and kinematic parameters in an estimator-in-the-loop manner. The upper level treats noise covariances and model parameters as optimization variables, while the lower level executes a full-information estimator. Differentiating through the estimator allows direct optimization of trajectory-level objectives, resulting in accurate and consistent state estimates. We validate our approach on quadrupedal and humanoid robots, demonstrating significantly improved estimation accuracy and uncertainty calibration compared to hand-tuned baselines. Our method unifies state estimation, sensor, and kinematics calibration into a principled, data-driven framework applicable across diverse robotic platforms.
I Introduction
State estimation is fundamental to the autonomy of robotic systems, providing the basis for planning and control. Robots must infer their states–such as position, velocity, and orientation-from noisy sensor data, often under challenging real-world conditions. Classical approaches such as the Kalman filter and its extensions [kalman1960, bai2023state] remain central, and have been successfully deployed in applications ranging from ground vehicles [thrun2005probabilistic] and aerial robots [mohamed2021momentum] to quadrupeds and humanoids [bloesch2013state, rotella2014state, he2025invariant, HLIP]. More recently, factor graph–based methods and smoothing techniques [kaess2012isam2, dellaert2017factor] have enabled large-scale estimation for simultaneous localization and mapping (SLAM) [cadena2017past] and visual–inertial odometry [mur2017orb, usenko2019visual].
A persistent challenge in these methods is the specification of noise covariance matrices. The process noise covariance captures uncertainty in the dynamics and actuation, while the measurement noise covariance characterizes sensor error models. In practice, both are difficult to obtain: manufacturer datasheets provide only partial information, while individual sensor calibration and physical system identification are both expensive and task-dependent. As a result, practitioners often rely on manual tuning [he2025invariant, hartley2020contact, kang24] - a heuristic, time-consuming process that can produce estimators that are either sub-optimal or inconsistent [thrun2005probabilistic] over out-of-distribution tasks. This problem is particularly acute for highly dynamic robots such as quadrupeds and humanoids, whose complex, high-dimensional hybrid dynamics and kinematics, and multi-modal sensor nonlinearities amplify the challenges to obtain precise state estimates [rotella2014state, hartley2020contact, he2025invariant, HLIP, kang24], especially under kinematic uncertainties.
Figure 1: Overview of the work with its application to a quadrupedal robot.
Several lines of work have been proposed to address this issue. Adaptive filtering methods [9044358] adjust covariances online based on innovation statistics, while expectation–maximization (EM) frameworks [em_mhe_2017, 10472604, yap2008simultaneous] iteratively refine noise parameters to maximize likelihood [6630924, 6482657]. Learning sensor parameters or noises parameters [wong2020variational, qadri2024learning, yap2008simultaneous] is another common approach. However, these methods often either assume no availability of ground truth data or only focus on decoupled calibrations. More importantly, their treatments have not been examined on modern legged robots that have multimodal sensor suites and complex dynamics.
In this work, we propose a bi-level optimization framework for simultaneous covariance and kinematics calibration for robotic state estimation. At the upper level, we treat covariance matrices and uncertain kinematics as optimization variables. At the lower level, we solve a full-information estimator using the Maximum a Posteriori (MAP) formulation given the current covariance estimates and kinematics information. By differentiating through the estimator [amos2017optnet, icml2022implicit], we directly optimize parameters to be calibrated with respect to trajectory-level objectives of minimizing the errors between the estimates and the measurable ground truth. This approach eliminates manual tuning, enforces physical consistency through structured parameterization and constraints, and generalizes across sensor modalities and robots.
We validate our approach on quadrupedal and bipedal robots, comparing against hand-tuned baselines. Across both platforms, our method calibrate the parameters to physically plausible values and significantly reduces estimation error and improves uncertainty consistency. These results highlight bi-level optimization for calibration as a principled and practical tool for bridging the gap between theoretical estimation frameworks and the demands of real-world robotics.
II Related Work
Legged Robotic State Estimation: State estimation for legged robots has been extensively studied due to the challenges posed by intermittent ground contacts, actuator uncertainties, and highly dynamic motions. Early approaches utilize Extended Kalman Filters (EKFs) to fuse proprioceptive sensing - Inertial Measurement Unit (IMU), contact sensors, and joint encoders - with kinematic constraints to achieve drift-reduced estimates on quadrupedal [bloesch2013state], bipedal [teng2021legged] and humanoid robots [rotella2014state]. Recent focuses of using EKFs on legged robots employ the invariant properties of the SE(3) manifold, and thus yield more accurate estimates [he2025invariant, hartley2020contact]. Another line of work leveraged factor graphs [dellaert2017factor, yang2024real] and smoothing or windowed formulations such as Moving Horizon Estimation (MHE) [kang24] to integrate proprioceptive and exteroceptive sensing data at scale. Despite these advances, most approaches rely on fixed or manually tuned noise covariances, limiting the accuracy of estimates, especially in the presence of kinematic uncertainty. Our work addresses this gap by introducing a bi-level optimization framework that calibrates covariance matrices and uncertain kinematics simultaneously in a principled manner.
Sensor Calibration: Sensor calibration and noise characterization have also been studied extensively in robotics, since the accuracy of estimation strongly depends on the quality of sensor models. Traditional approaches focus on extrinsic and intrinsic calibration of individual sensors, such as cameras [rehder2016extending], LiDARs [li2023automatic], and IMUs [fu2021high]. At the theoretical level, several works address the identification of noise statistics. For instance, expectation–maximization (EM) algorithms iteratively refine noise models to maximize likelihood [em_mhe_2017, 10472604, yap2008simultaneous]. Recent efforts explored data-driven estimation of uncertainty, including machine learning to optimize covariances from raw sensory inputs [qadri2024learning, wang2023neural]. Despite these advances, most existing techniques commonly are tailored to specific sensors rather than complex legged robots; they often assume restrictive statistical models, or require careful initialization. In contrast, our bi-level optimization jointly calibrates both process and measurement noise with uncertain kinematics across diverse and complex legged robotic platforms.
System Identification: System identification (SysID) has long been a cornerstone of robotics, enabling accurate modeling of dynamics, kinematics, and actuation. Classical techniques for parameter estimation of rigid-body kinematics or dynamics exploit regressions of the geometric or Lagrangian models [leboutet2021inertial, 7987066, khorshidi2025physically] subject to physical constraints. Gaussian processes or neural networks approximate residuals not captured by physics-based models [hwangbo2019learning] are commonly used in calibration. Recent works [yang2022online, burgul2024online] focus on online kinematics calibration via state estimators. Despite significant progress, SysID typically treats noise statistics, model parameters, and estimator design separately. Our bi-level optimization approach bridges this gap by jointly calibrating positive definite covariance matrices and kinematic parameters, highlighting the natural synergy between system identification and state estimation.
III Preliminaries
We begin by introducing the common modeling methods and techniques to address the state estimation of legged robots. The critical states to be estimated rather than measured are the linear velocity and orientation of their torso. We assume the robot is equipped with common proprioceptive sensors: joint encoders, IMU, and contact sensing/detection on the feet. Without loss of generality, we do not include additional exteroceptive sensors such as LiDAR or Cameras; yet, our method does not necessarily exclude fusing the exteroceptive sensing with the proprioceptive sensors.
We first introduce the models of legged state estimation that are widely used in the literature [rotella2014state, bloesch2013state, hartley2020contact, kang24, he2025invariant]. Then, we will elaborate the preliminaries of the mathematics on formulating state estimation as an optimization problem that we will employ in our calibration work. Throughout, $\mathrm{W}$ denotes the world frame and $\mathrm{B}$ the body (base) frame.
Figure 2: Illustration of the common sensing capabilities on legged robots: Quadrupedal Go1 and Humanoid G1 from Unitree Robotics.
III-A Modeling of Legged Robot Estimation
Legged robot state estimation is generally formulated by a nonlinear process and measurement model that jointly describe the system dynamics and sensor observations.
III-A1 Process Model
We consider a discrete-time nonlinear process model for state $\mathbf{x}_{k}$ and input $\mathbf{u}_{k}$ at time step $k=0,\ldots,T-1$ ($T$ denotes the length of the discrete horizon):
$\mathbf{x}_{k+1}=\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{w}_{k},\quad\mathbf{w}_{k}\sim\mathcal{N}(0,Q_{k}).$ (1)
The robot-centric state stacks the pose and the velocity of the floating-base, foot positions, and IMU biases: $\mathbf{x}_{k}=\begin{bmatrix}p^{\top}&v^{\top}&q^{\top}&p_{\text{foot},k}^{\top}&b_{a}^{\top}&b_{\omega}^{\top}\end{bmatrix}^{\top}$, where $p$ and $v$ are the linear position and velocity of the torso in $\mathrm{W}$, $q$ denotes the unit quaternion that parameterizes the rotation matrix $\mathbf{R}_{\mathrm{WB}}\in SO(3)$ of the torso w.r.t. $\mathrm{W}$, $p_{\text{foot},k}$ contains all the foot positions that are described in $\mathrm{W}$, and $b_{a},b_{\omega}\in\mathbb{R}^{3}$ are the biases terms of the IMU in $\mathrm{B}$. For later use, we write the state as a product-manifold element $\mathbf{x}_{k}=(\mathbf{T}_{k},\mathbf{e}_{k},b_{a,k},b_{\omega,k})\in\mathrm{SE}(3)\times\mathbb{R}^{n_{\mathrm{eucl}}}$, where $\mathbf{T}_{k}\in\mathrm{SE}(3)$ encodes the floating-base transformation with rotation parameterized by quaternion $q_{k}$, and $\mathbf{e}_{k},$ stacks the Euclidean sub-states, i.e., the base velocity $v_{k}$ and foot positions, $n_{\mathrm{eucl}}$ denotes the dimension of the Euclidean substate $\mathbf{e}_{k}$. The IMU measures the body acceleration $a$ and angular velocity $\omega$ with the standard additive biases and noises:
$\tilde{a}$ $=a+b_{a}+\delta_{a},$ $\delta_{a}$ $\sim\mathcal{N}(0,Q_{a}),$ (2)
$\tilde{\omega}$ $=\omega+b_{\omega}+\delta_{\omega},\quad$ $\delta_{\omega}$ $\sim\mathcal{N}(0,Q_{\omega}),$
and the biases have random walk dynamics:
$b_{a}^{+}$ $=b_{a}+\delta_{b_{a}},$ $\delta_{b_{a}}$ $\sim\mathcal{N}(0,Q_{b_{a}}),$ (3)
$b_{\omega}^{+}$ $=b_{\omega}+\delta_{b_{\omega}},$ $\delta_{b_{\omega}}$ $\sim\mathcal{N}(0,Q_{b_{\omega}}).$
Using these IMU measurements for state propagation, the discrete-time updates in (1) are:
$p^{+}$ $=p+v\,\Delta t+\tfrac{1}{2}\!\left(\mathbf{R}_{\mathrm{WB}}(\tilde{a}-b_{a})+g\right)\!\Delta t^{2}\;+\;\delta_{p},$
$v^{+}$ $=v+\left(\mathbf{R}_{\mathrm{WB}}(\tilde{a}-b_{a})+\mathbf{g}\right)\!\Delta t\;+\;\delta_{v},$
$q^{+}$ $=\zeta\!\left((\tilde{\omega}-b_{\omega})\,\Delta t\right)\otimes q\;+\;\delta_{q},$
$p_{\text{foot}}^{+}$ $=p_{\text{foot}}+\delta_{\text{foot}},\quad b_{a}^{+}=b_{a}+\delta_{b_{a}},\quad b_{\omega}^{+}=b_{\omega}+\delta_{b_{\omega}},$
where $\otimes$ denotes quaternion multiplication and $\zeta(\cdot)$ maps a small rotation vector to the corresponding unit quaternion, $g$ is gravitational acceleration in $\mathrm{W}$. The process noises $\delta_{(\cdot)}$ are modeled as zero-mean Gaussian $\mathcal{N}(0,Q_{(\cdot)})$, which we collect into a stacked random vector: $\mathbf{w}\!=\![\delta_{p}^{\top},\delta^{\top}_{v},\delta^{\top}_{q},\delta^{\top}_{\text{foot}},\delta^{\top}_{b_{a}},\delta^{\top}_{b_{\omega}}]^{\top}\!\sim\mathcal{N}\!\big(0,Q_{\mathbf{w}}\big)$, and $Q_{\mathbf{w}}=\operatorname{blkdiag}\!\big(Q_{p},Q_{a},Q_{\omega},Q_{\mathrm{foot}},Q_{b_{a}},Q_{b_{\omega}}\big)$. In practice, the random-walk covariances are scaled proportionally to the sampling period $\Delta t$.
III-A2 Measurement Model
The multi-sensor measurements $\mathbf{y}_{k}$ are functions of the state and input with Gaussian noise:
$\mathbf{y}_{k}=\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{n}_{k},\qquad\mathbf{n}_{k}\sim\mathcal{N}(0,R_{k}).$ (4)
The observation model $\mathbf{h}(\cdot)$ aggregates all available sensors. Joint encoders provide the measurements of the joint angles $\tilde{\alpha}$ and angular velocities $\dot{\tilde{\alpha}}$, which are assumed with Gaussian noises: $\tilde{\alpha}=\alpha+\delta_{\alpha}$ and $\dot{\tilde{\alpha}}=\dot{\alpha}+\delta_{\dot{\alpha}}$, where $\delta_{\alpha}\!\sim\!\mathcal{N}(0,R_{\alpha})$ and $\delta_{\dot{\alpha}}\!\sim\!\mathcal{N}(0,R_{\dot{\alpha}})$. Given leg kinematics, forward kinematics $fk(\cdot)$ and Jacobian $J(\cdot)$ yield the measurements of the foot positions and velocities w.r.t. the torso:
[…]
$y_{v}$ $=v+\delta_{y_{v}},$
[…]
where $\cdot$$_{\times}$ denotes skew-symmetric map, $\delta_{\mathrm{pf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{pf}})$ and $\delta_{\mathrm{vf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{vf}})$ encapsulates kinematics errors and encoder/gyroscope noises [teng2021legged]. Here $\tilde{(\cdot)}$ denotes a raw sensor measurement.
For simplicity, here we consider a single-leg situation; the multi-leg case follows by block-diagonal stacking. Let the proprioceptive noise $\delta z\!=\![\delta\alpha^{\top}\ \delta\dot{\alpha}^{\top}\ \delta\omega^{\top}]^{\top}$, and thus,
$\delta z\sim\mathcal{N}\!\left(0,\;R_{z}\right),R_{z}=\mathrm{blkdiag}\big(R_{\alpha},R_{\dot{\alpha}},Q_{\omega}\big)\in\mathbb{S}_{++}^{9},$ (6)
where the operator $\mathrm{blkdiag}(\cdot)$ forms a block-diagonal matrix from its arguments and $\mathbb{S}_{++}^{n}$ denotes the cone of $n\times n$ symmetric positive-definite matrices.. A first-order perturbation of the foot channels in (5) gives $\delta y_{\mathrm{pf}}\;=\;G_{p}\,\delta z$, $\delta y_{\mathrm{vf}}\;=\;G_{v}\delta z$, with Jacobian matrices as:
$\begin{bmatrix}G_{p}\\ G_{v}\end{bmatrix}\!=\!\begin{bmatrix}J(\alpha)&\mathbf{0}_{3\times 3}&\mathbf{0}_{3\times 3}\\ \dot{J}(\alpha,\dot{\alpha})+\omega^{\times}J(\alpha)&J(\alpha)&-[\,fk(\alpha)\,]_{\times}\end{bmatrix}\in\mathbb{R}^{6\times 9}.$
Thus, the induced measurement-noise covariances are $R_{\mathrm{pf}}\;=\;G_{p}\,R_{z}\,G_{p}^{\top}$, $R_{\mathrm{vf}}\;=\;G_{v}\,R_{z}\,G_{v}^{\top}$. If residuals terms $\mathbf{n}$ (8) are formed in $\mathrm{W}$ via $\mathbf{n}_{\mathrm{W}}=\mathbf{R}_{\mathrm{WB}}\;\mathbf{n}_{\mathrm{B}}$, the corresponding covariances becomes $R\mapsto\mathbf{R}_{\mathrm{WB}}\,R\,\mathbf{R}_{\mathrm{WB}}^{\top}$.
A binary contact flag $c\in\{0,1\}$, derived from pressure- or proprioceptive-based contact sensing, modulates the leg-odometry velocity channel: the base-velocity measurement is used only in stance ($c=1$) and suppressed in swing ($c=0$). In practice, the velocity channel’s contribution to the negative log-likelihood (objective) (III-B) (8) is multiplied by $c$.
III-B State Estimation as an MAP Problem
The optimal state estimator solves the Maximum A Posterior (MAP) problem:
$\hat{\mathbf{x}}^{*}$ $=\underset{\mathbf{x}_{[0,T]}}{\text{argmax}}\ \;{\mathbf{p}(\mathbf{x}_{[0,T]}|\mathbf{y}_{[0,T]})},$
$=\underset{\mathbf{x}_{[0,T]}}{\text{argmin}}\;\left(-\ln(\mathbf{p}(\mathbf{y}|\mathbf{x}))-\ln(\mathbf{p}(\mathbf{x}))\right),$ (7)
where the notation $(\cdot)_{[i,j]}$ denotes the sequence of vectors that starts from time index $i$ and ends at time index $j$. $\mathbf{p}(\mathbf{x}_{[0,T]}|\mathbf{y}_{[0,T]})$ is the posterior probability of the state trajectory $\mathbf{x}_{[0,T]}$, and $\hat{(\cdot)}^{*}$ denotes the optimal estimated solution.
For the estimation problem with the above-mentioned process and measurement models with additive Gaussian noises, the MAP problem can be solved via the Full information Estimation (FIE) formulation, with model descriptions formulated as state constraints:
$\min_{\mathcal{X}_{[0,T]}}\ \Gamma(\mathbf{x}_{0})$ $+\sum_{k=0}^{T-1}{\mathbf{w}_{k}^{\top}Q_{k}^{-1}\mathbf{w}_{k}}+\sum_{k=0}^{T}{\mathbf{n}_{k}^{\top}R_{k}^{-1}\mathbf{n}_{k}},$ (8)
$\text{s.t.}\quad\mathbf{x}_{k+1}$ $=\mathbf{f}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{w}_{k},\ \forall k=0,\ldots,T-1,$ (9)
$\mathbf{y}_{k}$ $=\mathbf{h}(\mathbf{x}_{k},\mathbf{u}_{k})+\mathbf{n}_{k},\ \forall k=0,\ldots,T,$ (10)
$g(\mathbf{x}_{k})$ $\leq 0,\ \forall k=0,\ldots,T.$ (11)
where $\mathcal{X}_{[0,T]}$ is the state and noise trajectories up to $T$:
$\mathcal{X}_{[0,T]}$ $=\begin{bmatrix}\mathbf{x}_{[0,T]}^{\top}\;\mathbf{w}_{[0,T-1]}^{\top}\;\mathbf{n}_{[0,T]}^{\top}\end{bmatrix}^{\top}$,
$\mathbf{w}_{[0,T-1]}$ $=\begin{bmatrix}\mathbf{w}_{0}^{\top}&\mathbf{w}_{1}^{\top}&\cdots&\mathbf{w}_{T-1}^{\top}\end{bmatrix}^{\top}$,
$\mathbf{n}_{[0,T]}$ $=\begin{bmatrix}\mathbf{n}_{0}^{\top}&\mathbf{n}_{1}^{\top}&\cdots&\mathbf{n}_{T}^{\top}\end{bmatrix}^{\top}$.
$g(\cdot)$ denotes general state constraints. Given a Gaussian-distributed prior $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{x}_{\text{prior}},P_{0})$, the prior cost $\Gamma(x_{0})$ is defined as: $\Gamma(\mathbf{x}_{0})=(\mathbf{x}_{0}-\mathbf{x}_{\text{prior}})^{\top}P_{0}^{-1}(\mathbf{x}_{0}-\mathbf{x}_{\text{prior}})$.
The major drawback of FIE is that its computational cost grows with the horizon $T$, making online use practically intractable. MHE [kang24] instead uses a fixed past window of measurements and repeatedly solves the problem on a shifting horizon (as in MPC), yielding bounded and predictable cost. Because legged estimation is real-time, EKF variants with analytical updates are commonly used instead of windowed optimization. For offline calibration, we exploit all available data and therefore adopt FIE.
IV Problem Statement
With properly selections of the noise covariance matrices $Q$ and $R$, the commonly applied estimators such as Extended Kalman Filters (EKFs) [bloesch2013state, rotella2014state], invariance EKFs [hartley2020contact, he2025invariant], and MHE [kang24] can yield satisfying estimates. However, it is widely known that extensive expert tuning of $Q$ and $R$ is inevitable on complex robots, and it can easily lead to sub-optimal, inconsistent or even diverging estimates on real robots. Additionally, inaccurate kinematics can further degrade estimate stability and precision. Therefore, we aim to provide an automatic and tuning-free method to optimally calibrate $Q$ and $R$ along with uncertain kinematics so that any reasonable choices of the estimators, regardless of EKFs or MHEs depending on the computational budget, can yield precise and consistent estimation on diverse set of tasks.
IV-A Assumptions and Objective
Sensors and Ground Truth Measurements: We assume the robot is equipped with common proprioceptive sensors that are specified in previous section. The robot has a functional controller that allows it to locomote in an environment that has motion capture devices to provide accurate measurements of the torso pose and velocities, which will be used as ground truth, denoted at $\mathbf{x}_{\text{GT}}$. Note that $\mathbf{x}_{\text{GT}}$ does not cover all the states such as the IMU biases. We assume that the robot is controlled to realize certain tasks with which all the sensor information $\mathbf{Y}:=\{\mathbf{y}_{k}\}_{k=0}^{T}$ along with the ground truth measurements $\mathbf{x}_{\text{GT}}$ are provided to us.
Objectives: Our objective is thus to calibrate the optimal covariance matrices $Q$ and $R$ along with uncertain kinematic parameters so that the estimated results will be as close as possible to the ground truth using our process and measurement models. Without losing generality, for legged robots, we assume the last link on each leg has uncertain kinematic information; other linkage can be assumed to have uncertainties as well. This kinematic uncertainty on the last link (shin/calf) results in an uncertain foot position. Additionally, the motion capture device will capture markers that are placed on the torso. Therefore, besides uncertain shin kinematics, $\mathbf{x}_{\text{GT}}$ will not directly measure the robot state but with a constant kinematic offset, which will need to be calibrated along with the calibration of $Q$ and $R$.
IV-B Bi-level Optimization
We propose a bi-level optimization to solve the calibration problem. The optimization is given by:
$\underset{\bm{\theta}\in\mathbb{R}^{m}}{\min}\quad$ $\mathcal{L}\bigl(\hat{\mathbf{x}}\left(\bm{\theta}\right),\mathbf{x}_{\mathrm{GT}},\bm{\theta}\bigr),$ (12)
$\mathrm{s}.\mathrm{t.}\quad$ $\bm{\theta}\in\mathcal{C},$
$\hat{\mathbf{x}}\left(\bm{\theta}\right)=\mathop{\mathrm{arg}\,\min}\limits_{\mathbf{x}\in\mathbb{R}^{n\times(T+1)}}\;\mathcal{J}\bigl(\mathbf{x},\bm{\theta}\bigr).$
At the upper level, we minimize a differentiable loss $\mathcal{L}$ that measures the deviation between the estimated trajectory $\hat{\mathbf{x}}$ and the ground truth trajectory $\mathbf{x}_{\mathrm{GT}}$. We decompose the decision vector into covariance parts and geometric calibration parts $\bm{\theta}=\bigl(\bm{\theta}_{\mathrm{cov}},\bm{\theta}_{\mathrm{foot}},\bm{\theta}_{\mathrm{base}}\bigr)$, where $\bm{\theta}_{\mathrm{cov}}$ parameterizes the covariance matrices, $\bm{\theta}_{\mathrm{foot}}$ stacks the offsets on the foot kinematics, and $\bm{\theta}_{\mathrm{base}}\in\mathbb{R}^{3}$ is the constant offset from the base frame of the torso $\mathrm{B}$ to the motion-captured body frame rigidly attached to the torso $\mathrm{M}$ , expressed in $\mathrm{B}$. In particular, each $\bm{\theta}_{\mathrm{foot},j}\in\mathbb{R}^{3}$ is a constant vector expressed in the calf frame of leg $j$ $C_{j}$ that points from the foot to the contact point on the real robot. Thus, the actual foot position is:
$p^{\mathrm{W}}_{\mathrm{real},j}\;=\;p^{\mathrm{W}}_{{\mathrm{foot}},j}\;+\;\mathbf{R}_{\mathrm{WC_{j}}}\,\bm{\theta}_{\mathrm{foot},j},$ (13)
in which the offset $\bm{\theta}_{\mathrm{foot},j}$ is rotated from $C_{j}$ to the world frame by $\mathbf{R}_{\mathrm{WC}_{j}}$ and added to the nominal foot position $p^{\mathrm{W}}_{\mathrm{foot},j}$ to yield the actual contact location $p^{\mathrm{W}}_{\mathrm{real},j}$. This time-invariant offset represents kinematic errors that come from mounting tolerances or manufacture inaccuracy. Note that, the observation vector $\mathbf{y}(\bm{\theta})$ (10) and the kinematic mapping matrices $\mathbf{G}(\bm{\theta})$ (III-A2) depend explicitly on $\bm{\theta}_{\mathrm{foot}}$.
At the lower level, we fix the vector $\bm{\theta}$ and solve the MAP problem, where $\mathcal{J}$ is the MAP cost in (III-B), (8). Given a covariance specification $\Sigma(\bm{\theta})$, where $\Sigma(\bm{\theta})$ is the collection of $Q$ and $R$, the solution of the MAP yields an estimated trajectory $\hat{\mathbf{x}}(\bm{\theta})\in\mathcal{X}$.
The estimated state at time $k$ in trajectory is defined as $\hat{\mathbf{x}}_{k}\;=\;(\mathbf{T}_{k},\,\mathbf{e}_{k})\;\in\;\mathrm{SE}(3)\times\mathbb{R}^{n_{\mathrm{eucl}}}$, where $\mathbf{T}_{k}$ denotes the base pose and $\mathbf{e}_{k}$ collects Euclidean sub-states such as base velocity and contact foot positions. The full trajectory lies on the manifold $\mathbf{X}\;\in\;\mathrm{SE}(3)^{T+1}\times\mathbb{R}^{n_{\mathrm{eucl}}(T+1)}$.
IV-C Constraint Set $\mathcal{C}$
Given a positive-definite matrix $\Sigma_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$, we define its vectorization $\bm{\theta}_{\text{cov}}=\operatorname{vec}(\Sigma_{i})\in\mathbb{R}^{d_{i}(d_{i}+1)/2}$ by stacking the independent entries (upper triangular part including the diagonal) in a fixed order. During optimization, the parameter vector must remain in the constraint set $\mathcal{C}=\left\{\bm{\theta}\in\mathbb{R}^{m}\,\middle|\,\underline{\bm{\theta}}\leq\bm{\theta}\leq\overline{\bm{\theta}},\quad\Sigma_{i}(\bm{\theta})\in\mathbb{S}^{d_{i}}_{++},\;\forall i\right\}$, where $\underline{\bm{\theta}},\overline{\bm{\theta}}\in\mathbb{R}^{m}$ denote element-wise lower and upper bounds on each parameter, including both variances, correlations, and kinematic paramters. The set $\mathbb{S}^{d_{i}}_{++}$ denotes the cone of symmetric positive definite matrices of size $d_{i}\times d_{i}$.
To ensure statistical and physical feasibility, the constraint set $\mathcal{C}$ imposes two requirements. First, each covariance block $\Sigma_{i}$ is constrained to lie in $\mathbb{S}^{d_{i}}_{++}$ by the property of covariance matrices. Second, all parameter entries are bounded within predefined intervals $\underline{\bm{\theta}},\overline{\bm{\theta}}]$, reflecting prior sensor knowledge in the form of individual bounds on variances and correlations. These bounds also serve to improve numerical stability by avoiding ill-conditioned solutions during optimization.
The convexity and compactness of the set $\mathcal{C}$: The cone $\mathbb{S}^{d_{i}}_{++}$ is convex but open cone. By imposing element-wise box constraints $\underline{\bm{\theta}},\overline{\bm{\theta}}]$, we obtain a closed and bounded subset $\mathcal{C}$, which remains convex and is therefore compact.
V Method
We now present our method to solve the bi-level optimization problem in (12). We adopt an iterative optimization approach: the lower-level problem (8) is solved using a standard interior-point method [wachter2006implementation], while the upper-level problem leverages its solution and Karush–Kuhn–Tucker (KKT) condition to update the parameters $\bm{\theta}$ via an adaptive Frank–Wolfe algorithm that iteratively refines the estimate until convergence. The index $t\in\mathbb{N}$ denotes the Frank–Wolfe iteration number.
Algorithm 1 Frank–Wolfe Algorithm
1:Input: $\mathcal{L}$, $\bm{\theta}^{t}$, $\mathcal{C}_{t}$, $t$
2:Gradient compute: $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})$ $\triangleright$ as in Sec. (V-B2)
3:Direction finding (LMO):
$\mathbf{s}^{\star}=\mathrm{argmin}\;$ $\;\mathbf{s}^{\top}\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})$
subject to $\;\mathbf{s}\in\mathcal{C}_{t}\subset\mathbb{R}^{m}$
4:Step-size: $\gamma_{t}\!\in\!(0,1]\leftarrow\textsc{LineSearch}\!\left(\mathcal{L},\nabla\mathcal{L},\bm{\theta}^{t},\mathbf{s}^{\star}\right)$
5:Update: $\bm{\theta}^{t+1}\leftarrow(1-\gamma_{t})\,\bm{\theta}^{t}+\gamma_{t}\,\mathbf{s}^{t}$
V-A Frank–Wolfe Algorithm for the Upper Level Optimization
Given that the constraint set $\mathcal{C}$ is convex and compact, we minimize the upper-level objective over $\mathcal{C}$ using the Frank-Wolfe algorithm, which preserves constraints without requiring projections. Given the upper-level gradient $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})$ computed as in Sec. (V-B2), each iteration solves a Linear Minimization Oracle (LMO) that minimizes the first-order approximation of $\mathcal{L}$ over $\mathcal{C}$, and updates the candidate solution by taking a convex combination of the current point with the LMO solution through line search.
To enhance numerical stability, each Frank–Wolfe subproblem is restricted to a local trust region around the current iterate $\bm{\theta}^{(t)}$: $\mathcal{C}_{t}=\mathcal{C}\cap\Bigl\{\bm{\theta}\in\mathbb{R}^{m}\;\big|\;\|\bm{\theta}-\bm{\theta}^{(t)}\|_{\infty}\leq\Delta_{t}\Bigr\}$, where $\Delta_{t}>0$ denotes the trust-region radius. This intersection preserves convexity and compactness, ensuring that the linear minimization oracle remains well-posed.
Linear Minimization Oracle: The LMO is implemented using two approaches:
(i) Constraint-based approach: As $\mathcal{L}$ is nonlinear in $\bm{\theta}$, in each Frank–Wolfe iteration we linearizes the objective at $\bm{\theta}^{t}$, which turns the direction-finding step into the optimization of a linear function over the intersection of the positive semidefinite cone with box constraint. This can be efficiently solved by standard SDP solvers. Concretely, we embed $\mathbf{s}$ into symmetric block matrices via fixed symmetric basis matrices $\{F_{i}\}$ and enforce strict positive definiteness:
$\min_{\mathbf{s}}$ $\quad\mathbf{s}^{\top}\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t}),$
s.t. $\quad\underline{\mathbf{s}}\leq\mathbf{s}\leq\overline{\mathbf{s}},\;\;\sum_{i}s_{i}F_{i}\;\succeq\;\epsilon I.$
(ii) Parameterization-based approach: Alternatively, we parameterize each covariance block as $\Sigma_{i}=L_{i}L_{i}^{\top}$, where $L_{i}$ stands for the lower triangular matrix with positive entries. This enforces $\Sigma_{i}$ being positive definite. In this approach, $\mathcal{C}$ reduces to box bounds on the free entries of $\{L_{i}\}$, and the LMO becomes a lightweight box-constrained linear program in those entries.
Armijo Line Search: In practice, predefined decreasing step-size schedules often induce oscillations in $\mathcal{L}$ and make $\|\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})\|_{2}$ converge only after many iterations. We therefore adopt Algorithm 2 to choose an adaptive step size $\gamma_{t}$ that yields the sufficient decrease per iteration along the feasible FW direction returned by the LMO. We select $\gamma_{t}$ to satisfy
$\mathcal{L}\!\left(\bm{\theta}^{t}+\gamma_{t}\big(\mathbf{s}^{\star}-\bm{\theta}^{t}\big)\right)\leq\mathcal{L}(\bm{\theta}^{t})\;+\;\rho\,\gamma_{t}\;\big(\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})\big)^{\top}\big(\mathbf{s}^{\star}-\bm{\theta}^{t}\big),$
where $\rho\in(0,1)$ enforces a fixed fraction of the predicted decrease. The accepted $\gamma_{t}$ from Algorithm 2 is then used to update $\bm{\theta}$ in Algorithm 1.
Algorithm 2 Armijo Line Search (backtracking for $\gamma_{t}$)
1:Parameters: $\rho\in(0,1)$, $\beta\in(0,1)$, $\gamma\leftarrow 1$
2:while
$\mathcal{L}\!\left(\bm{\theta}^{t}+\gamma(\mathbf{s}^{\star}\!-\bm{\theta}^{t})\right)>\mathcal{L}(\bm{\theta}^{t})+\rho\,\gamma_{t}\big(\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{t})\big)^{\top}\!\big(\mathbf{s}^{\star}\!-\bm{\theta}^{t}\big)$
do $\gamma\leftarrow\beta\,\gamma$
3:end while
4:Return: $\gamma_{t}\leftarrow\gamma$
V-B Cost Definition and Gradient Computation
Given that the full trajectory lies on the Lie-group manifold $\mathcal{M}=\mathrm{SE}(3)^{T+1}\times\mathbb{R}^{\,n_{\mathrm{eucl}}(T+1)}$ with $\mathbf{X}\in\mathcal{M}$, we denote the per-time-step state manifold by $\mathcal{G}=\mathrm{SE}(3)\times\mathbb{R}^{\,n_{\mathrm{eucl}}}$ with $\mathbf{x}\in\mathcal{G}$, which implies that $\mathcal{M}=\mathcal{G}^{T+1}$. The smooth operators on the manifold $\mathcal{G}$ are defined as $\boxplus:\ \mathcal{G}\times\mathbb{R}^{6+n_{\mathrm{eucl}}}\to\mathcal{G},\quad\boxminus:\ \mathcal{G}\times\mathcal{G}\to\mathbb{R}^{6+n_{\mathrm{eucl}}}$. We write $\log\!:\!\mathrm{SE}(3)\!\to\!\mathfrak{se}(3)$ and $\exp\!:\!\mathfrak{se}(3)\!\to\!\mathrm{SE}(3)$ for the Lie logarithm and exponential maps.
V-B1 Upper level loss function $\mathcal{L}$
For each time step in the state trajectory, the upper-level loss function $\mathcal{L}:\ \mathcal{G}\to\mathbb{R}$ is defined by measuring trajectory deviation through residual operator that respects the product-manifold geometry. For the $\mathrm{SE}(3)$ component, we use the geodesic distance induced by the left-invariant logarithmic map, while the Euclidean component relies on the Euclidean norm. The cost for the entire trajectory from time index 0 to $T$ is defined as:
$\mathcal{L}$ $\big(\hat{\mathbf{x}},\mathbf{x}_{\mathrm{GT}}\big)=\frac{1}{2}\sum_{k=0}^{T}\big\|\hat{\mathbf{x}}_{k}\boxminus\mathbf{x}_{\mathrm{GT},k}\big\|_{2}^{2},$ (14)
$=\frac{1}{2}\sum_{k=0}^{T}\left(\big\|\left(\log\big(\mathbf{T}_{k}\,\mathbf{T}_{\mathrm{GT},k}^{-1}\big)\right)^{\vee}\big\|_{2}^{2}+\big\|\mathbf{e}_{k}-\mathbf{e}_{\mathrm{GT},k}\big\|_{2}^{2}\right),$
where $(\cdot)^{\vee}:\mathfrak{se}(3)\to\mathbb{R}^{6}$ maps a matrix in $\mathfrak{se}(3)$ to the twist.
V-B2 Gradient
Each Frank–Wolfe iteration in Sec. (V-A) requires the gradient w.r.t. the calibration parameters, i.e., $\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$, so we begin by formalizing gradients on the product manifold via left-invariant Lie-group Jacobians. The left Jacobian of $\mathcal{L}$ at $\hat{\mathbf{x}}\in\mathcal{G}$ is the linear map, which is:
$\frac{{}^{e}D\mathcal{L}(\hat{\mathbf{x}})}{D\hat{\mathbf{x}}}=\lim_{\epsilon\to 0}\frac{\mathcal{L}\!\big(\hat{\mathbf{x}}\boxplus(\epsilon\tau)\big)-\mathcal{L}(\hat{\mathbf{x}})}{\epsilon},$ (15)
for any $\tau\in T_{e}(\mathcal{G})\cong\mathbb{R}^{6+n_{\mathrm{eucl}}}$. For the $\mathrm{SE}(3)$ component, with canonical basis $\{\bm{\xi}_{i}\}\subset\mathfrak{se}(3)$,
$\left[\frac{{}^{e}D\mathcal{L}(\mathbf{T})}{D\mathbf{T}}\right]_{i}=\lim_{\epsilon\to 0}\frac{\mathcal{L}\big(\exp(\epsilon\bm{\xi}_{i})\cdot\mathbf{T}\big)-\mathcal{L}(\mathbf{T})}{\epsilon}.$ (16)
Similarly, the left Jacobian of the estimated solution $\hat{\mathbf{x}}(\bm{\theta})$ w.r.t. $\bm{\theta}$ is $^{e}\!D\hat{\mathbf{x}}(\bm{\theta})/D\bm{\theta}$ where is defined as
$\frac{{}^{e}D\hat{\mathbf{x}}(\bm{\theta})}{D\bm{\theta}}=\lim_{\epsilon\to 0}\frac{\hat{\mathbf{x}}(\bm{\theta}+\epsilon\nu)\;\boxminus\;\hat{\mathbf{x}}(\bm{\theta})}{\epsilon},$ (17)
for any $\nu\in\mathbb{R}^{m}$. For the $\mathrm{SE}(3)$ component,
$\left[\frac{{}^{e}D\mathbf{T}}{D\bm{\theta}}\right]_{j}=\lim_{\epsilon\to 0}\frac{\left(\log\!\big(\mathbf{T}\cdot\tilde{\mathbf{T}}_{j}^{-1}\big)\right)^{\vee}}{\epsilon},$ (18)
where $\tilde{\mathbf{T}}_{j}$ is the perturbed pose obtained from $\bm{\theta}\!\mapsto\!\bm{\theta}+\epsilon\,\mathbf{e}_{j}$.
Analytical Gradient Computation: The above-mentioned limit definitions allow finite-differencing for computation. However, with long horizons and high-dimensional upper-level optimization variables, the computational cost of numerical differentiation remains considerable even when parallelized on modern CPUs. Therefore, we employ analytical derivatives instead. The upper-level loss is aggregated over time via the product-manifold residual in (14). Using the chain rule on $\mathcal{M}=\mathcal{G}^{T+1}$,
$\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})\;=\;\left(\frac{D\mathcal{L}(\hat{\mathbf{x}}(\bm{\theta}))}{D\hat{\mathbf{x}}}\right)\left(\frac{D\hat{\mathbf{x}}(\bm{\theta})}{D\bm{\theta}}\right)\;+\;\frac{\partial\mathcal{L}}{\partial\bm{\theta}},$ (19)
where $D\mathcal{L}(\hat{\mathbf{x}})/D\hat{\mathbf{x}}\in\mathbb{R}^{1\times d}$ and $D\hat{\mathbf{x}}(\bm{\theta})/D\bm{\theta}\in\mathbb{R}^{d\times m}$ are formed and stacked over the horizon with $d\!=\!(6+n_{\mathrm{eucl}})(T\!+\!1)$. To differentiate through the lower-level estimator, the optimal solution $\hat{\mathbf{x}}^{\star}(\bm{\theta})$ is implicitly defined by the first-order optimality conditions of problem (8). We treat the noise variables $\mathbf{w}$, $\mathbf{n}$ and the multipliers $\bm{\lambda}$ as implicit functions of the state $\hat{\mathbf{x}}^{\star}$ and the parameter $\bm{\theta}$. The KKT residual is: $\mathcal{F}\!\left(\hat{\mathbf{x}}^{\star},\bm{\theta},\mathbf{y}(\bm{\theta}),\mathbf{G}(\bm{\theta})\right)=0$. Because of the kinematics offset parameters, the observation $\mathbf{y}$ (10) and the mapping matrices from joint kinematics to measurement space $\mathbf{G}\!=\!\{G_{i}\}$ (III-A2) depends explicitly on $\bm{\theta}_{\mathrm{foot},j}$ (13). By differentiating this system with respect to $\bm{\theta}$, we obtain the sensitivity of the optimal state trajectory by implicit function theorem:
$\frac{\partial\hat{\mathbf{x}}^{\star}}{\partial\bm{\theta}}=-\;\left(\frac{\partial\mathcal{F}}{\partial\hat{\mathbf{x}}^{\star}}\right)^{-1}\left(\frac{\partial\mathcal{F}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{y}}\frac{\partial\mathbf{y}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{G}}\frac{\partial\mathbf{G}}{\partial\bm{\theta}}\right),$
where the right-hand side collects explicit parameter derivatives, implicit measurement dependence, and the additional contribution through $\mathbf{G}$. This states that the total effect of $\bm{\theta}$ on the optimal inner variables is the sum of a direct term and two chain–rule terms passing through $\mathbf{y}(\bm{\theta})$ and $\mathbf{G}(\bm{\theta})$.
In our estimation formulation, the decision vector $\mathbf{x}$ stacks all states over the whole window of data. Because dynamics and measurements couple locally, the KKT system is large yet sparse with a banded or block-tridiagonal pattern. Therefore, instead of forming dense matrix inverses, we express
$\left(\frac{\partial\mathcal{F}}{\partial\hat{\mathbf{x}}^{\star}}\right)\,Z=-\left(\frac{\partial\mathcal{F}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{y}}\frac{\partial\mathbf{y}}{\partial\bm{\theta}}+\frac{\partial\mathcal{F}}{\partial\mathbf{G}}\frac{\partial\mathbf{G}}{\partial\bm{\theta}}\right),$
where we use sparse linear solvers to solve $Z:=(\partial\hat{\mathbf{x}}^{\star})/(\partial\bm{\theta})$ efficiently. This makes each application of the inverse a low incremental cost, achieved through back-substitutions with a high-performance solver such as PyPardiso.
VI Results
We evaluate our calibration approach on three robots: a bipedal robot STRIDE [kang24] and two quadrupedal robot Go1 and B1 from Unitree. The data used from STRIDE and Go1 are obtained from simulation in Matlab and MuJoCo environment in their open source repository [kang24]. We obtained the data of B1 on hardware in the motion capture room with 12 Opti-track Cameras, with a combination of Prime 13 and 22.
The bi-level optimization is implemented in Python and C++. The outer loop iteration as LMO is solved by MOSEK; the inner loop FIE is solved with IPOPT [wachter2006implementation]. The analytical gradients are generated via CasADi [Andersson2019] and Pinocchio [carpentier:hal-01866228].
We concisely focus on presenting three core results as follows. All the results are consistent across these robots, regardless of their morphology and dynamics, showing successful calibrations of noise covariances and kinematic parameters with significantly improved robot estimates.
Figure 3: The robot STRIDE (a), quadrupedal robot Go1 (b), and B1 (c) are used in the evaluation (Pictures are used with permission).
Joint Sensor Noise & Kinematics Calibration: For the robot STRIDE, we mainly evaluate the joint calibration on the sensor noises and robot kinematics since their dynamics processes are deterministic in Matlab. We inject sensor noises into the reading along with an articulated shin length offset to the estimator. Fig. 4 shows the calibration results. The loss successfully reduced through the iterations of the bi-level optimization, and the estimated velocity converges to the ground truth; the kinematic parameters converge as well.
Figure 4: Calibration results on STIRDE: (a) the convergence of the upper loss function and norm of the gradients, (b) the convergence of kinematics bias w.r.t. ground truth, (c) and (d) the linear velocities of the ground truth, initial estimates, and calibrated estimates.
Joint Process & Sensor Noise Calibration: For the robot Go1, we evaluate the calibration using data generated by MuJoCo’s time-stepping, multi-contact simulation. Sensor noise is defined in the MJCF/XML and evaluated at every step as a function of the simulated state. The loss and gradient norm successfully converge through the iterations of the bi-level optimization. Additionally, we inject an offset between the torso and ground truth to model motion-capture offset. As optimization proceeds, the estimated offset converges to the pre-defined ground-truth value.
Figure 5: Calibration results on Go1: convergences of cost and gradient (top), and kinematic offset (bottom).
Joint Noise & Kinematics Calibration: Last and most importantly, we calibrate the robot B1 using hardware data. The robot is controlled to walk in the motion-capture room. The torso motion is measured by the motion capture system, taken as ground truth with a constant unknown offset to the base frame position of the torso. The orientation measurement of the torso is assumed to be unbiased, because the torso and the marked rigid body are well-aligned in the beginning of the experiment. Additionally, we assume all the shins of B1 have kinematic errors from the factory URDF. Thus, besides the noise covariances, the kinematic errors to be calibrated are 15-dimensional. After we obtained the sensory data along with the ground truth measurements, we applied our calibration to B1. Fig. 7 shows the improved estimates of the torso’s orientation and linear velocity, and Fig. 6 shows the kinematics converge to reasonable values: in our experimental setup, the marked torso location is indeed about centimeters away from the base frame location. Table I shows the quantitative results in terms of RMSE; Fig. 8 shows the new evaluation on a different segment of data using calibrated covariances and kinematics, which still yields highly accurate estimates. To sum up, our joint calibration significantly improves the estimates by very large margins.
TABLE I: RMSEs of the estimates on B1 robot hardware.
Estimation Metric Before Calib. After Calib. New Eval.
$\text{RMSE}_{v}$ [m/s] 0.2658 0.0610 0.0706
$\text{RMSE}_{\text{Euler}}$ [rad] 0.3457 0.0151 0.0514
Figure 6: Convergence of the kinematics calibration on B1 robot hardware: (top) the foot offset of the rear right leg and (bottom) the base offset.
Figure 7: Calibration results on B1 hardware in terms of linear velocity and torso orientation.
Figure 8: Evaluation results on B1 hardware in terms of linear velocity and torso orientation.
VII Conclusion and Future Work
We presented a bi-level optimization framework for jointly calibrating noise covariances and kinematics in legged robot state estimation. By differentiating through a full-information estimator, our method removes manual tuning and improves accuracy and consistency across simulated and real robots. Future work will focus on reducing computational cost, analyzing observability, and extending the approach to dynamics and inertia identification for fully self-calibrating systems.